If the fundamental frequency of cyclotron emission is 2.0 x 1014 Hz and the angular velocity of particles is 4.0 x 1014 rad/sec-1, find its wavelength.
Answer #1

We know that the wavelength is

=C/f

Where c=velocity of light in free space

f= frequency . Now the wavelength =(3×10^8)/(2×10^14)

=1.5×10^-6 m